Circular No. 8268

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004B IN IC 390 Further to IAUC 8265, K. Itagaki reports (via S. Nakano) the discovery

of another possible of a supernova (mag 17.8) on numerous unfiltered CCD frames taken on Jan. 12.569 UT with a 0.60-m f/5.7 reflector. The new object is located at α = 4^{h}42^{m}04^{s}.71, δ = −7^{o}12′08′′.2 (equinox 2000.0), which is 12′′ east and 15′′ north of the center of IC 390. Itagaki found nothing at this position on his frames taken on 2003 Oct. 27 (limiting mag 18.3) or on the Digitized Sky Survey. Nakano also reports that a confirming unfiltered CCD image was obtained on Jan. 14.599 by R. Kushida in bad seeing conditions; Y. Kushida reports that SN 2004B then appeared at mag 17.9 at position end figures 04^{s}.74, 07′′.4. H. Yamaoka, Kyushu University, adds that nothing is visible at this position on a red Digitized Sky Survey image taken on 1995 Nov. 14.

SUPERNOVAE 2003ld, 2003lo, 2003lp, AND 2003lq T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ld (cf. IAUC 8260), obtained by M. Calkins on Jan. 13.16 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova; adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 4213 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6000 km/s. A spectrogram of SN 2003lo (cf. IAUC 8261), obtained on Jan. 13.18, shows it to be a type-IIn supernova; the spectrum shows a blue continuum with both broad and narrow hydrogen emission lines. A spectrogram of SN 2003lp (cf. IAUC 8261), obtained on Jan. 13.46, shows it to be a type-II supernova; adopting the NED recession velocity of 2702 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8200 km/s. A spectrogram (range 370-750 nm) of SN 2003lq (cf. IAUC 8262), obtained on Jan. 13.15, shows it to be a type-Ia supernova ∼ a week past maximum. Adopting the NED recession velocity of 7271 km/s for the host galaxy, the supernova expansion velocity is 9700 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 6 ± 2 days after maximum light.

COMET C/2002 T7 (LINEAR) Visual total-magnitude estimates with 80-mm binoculars: 2003 Dec.

2.45 UT, 9.3 (C. S. Morris, Fillmore, CA); 25.77, 8.1 (R. Haver, Frasso Sabino, Italy); 2004 Jan. 13.99, 8.3 (A. Amorim, Florianopolis, Brazil).

2004 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green